comcertain motor oil is packed in tins holding 5 litres each. The filling machine can maintain this but with a S.D. of 0.15 litre. Two samples of 36 tins each are taken from the production line. If the sample means are 5.20 and 4.95 litres respectively, can we be 99% sure that the sample have come from a population of 5 litres?
1
Expert's answer
2020-05-04T19:35:47-0400

"1) N_1=36, \overline{x}_1=5.2"

"H_0: a=a_0=5, H_1:a>a_0=5\text{ (one-sided)}\\
a \text{ is population mean}\\
\alpha=0.01\\
\sigma=0.15\\
\text{We assume that volume has normal distribution with mean } a\\
\text{and standard deviation } \sigma=0.15.\\
\text{We will use the following random variable:}\\
U=\frac{(\overline{X}-a_0)\sqrt{n}}{\sigma}\\
\overline{X}\text{ --- random value of sample mean}\\
u_{obs}=\frac{(5.2-5)\sqrt{36}}{0.15}=8\\
\Phi(u_{cr})=\frac{1-2\alpha}{2}=0.49\\
u_{cr}=2.33\\
\Phi(x)\text{ --- Laplace function}\\
(2.33,\infty)\text{ --- critical region}\\
u_{obs} \text{ is in the critical region. So we reject } H_0.\\
\text{We cannot be 99 percent sure that the sample have come from}\\
\text{a population of 5 litres}." "2) N_2=36, \overline{x}_2=4.95\\
H_0: a=a_0=5, H_1:a<a_0=5\text{ (one-sided)}\\
a \text{ is population mean}\\
\alpha=0.01\\
\sigma=0.15\\
\text{We assume that volume has normal distribution with mean } a\\
\text{and standard deviation } \sigma=0.15.\\
\text{We will use the following random variable:}\\
U=\frac{(\overline{X}-a_0)\sqrt{n}}{\sigma}\\
\overline{X}\text{ --- random value of sample mean}\\
u_{obs}=\frac{(4.95-5)\sqrt{36}}{0.15}=-2\\
(-\infty,-u_{cr})\text{ --- critical region where } u_{cr}=2.33\\
u_{obs} \text{ is not in the critical region. So we accept } H_0.\\
\text{We can be 99 percent sure that the sample have come from}\\
\text{a population of 5 litres}."